Grade 10Mechanics

a particle of mass m is acted upon by a force F=Focos{(pie/2)t}. the distance travelledby the particle when it stops first[initial velocity of the particle is zero) is .Is the answer 8f/m pi^2

Ramesh Tiwari
6 Years agoGrade 10

2 Answers

Arun
6 Years ago
P=FV
P=mav
P=mv . dv/ds . v
P/m.ds =v^2.dv
Integrate both sides under limits
V from v1 to v2
ds from 0 to S
We get
(v^3/3)v1tov2=P/m.s
1/3(v2^3-v1^3)=P/m.S
M/3P(v2^3-v1^3)=S
Vikas TU
6 Years ago
Dear student 
v = 2* f1 / (pi * m)
integrate with limits 0 to 1
as f=0 when t = 1
Good Luck …...........